-1/4(x-8)+1/2(x+2)=x+15
(-1/4*x)-(-1/4*8)+(1/2*x)+(1/2*2)=x+15
-1/4x-2+1/2x+1=x+15
3/4x-1=x+15
(3/4x-3/4)-1=(x-3/4)+15
x-1=1/4x+15
And now I'm completely lost. Thanks for any help you can give.

Answer by Alan3354(30993)

-1/4(x-8)+1/2(x+2)=x+15
(-1/4*x)-(-1/4*8)+(1/2*x)+(1/2*2)=x+15
-1/4x-2+1/2x+1=x+15 ************************-(-1/4*8) = +2
3/4x-1=x+15
----------------
-1/4(x-8)+1/2(x+2)=x+15
(-1/4*x)-(-1/4*8)+(1/2*x)+(1/2*2)=x+15
-1/4x + 2 + 1/2x + 1 = x+15
(1/4)x + 3 = x+15
(1/4)x = x + 15
x = 4x + 60
-3x = 60
x = -20